# Long Chain-of-Thought for Reasoning Large Language Models

### v3

@@ 6.2 Internal Exploration
As noted in Chu et al. [109], Shen et al. [527], and Yeo et al. [769], SFT serves as a memory process, while RL enhances generalization [287]. Specifically, SFT stabilizes the model’s output format, whereas RL improves its generalization capacity, which can increase learning efficiency by up to eight times in tasks such as mathematical reasoning [518]. Consequently, as shown in Figure 10, leading research emphasizes the role of RL and reward strategies in enhancing the exploration capabilities of LLMs without external assistance. The performance comparison is presented in Table 5.
@@ 6.2.1 RL Strategies
(1) Reward-model-free RL: The first series of work focuses on RL optimization algorithms. Additionally, OREO [621] propose an offline RL method that optimizes the soft Bellman equation, improving credit assignment for multi-step reasoning tasks and outperforming existing approaches in fields like mathematics and agent control. Liu et al. [380] propose Direct Advantage Policy Optimization (DAPO), a novel offline RL method that leverages a separately trained critic to evaluate the accuracy of each reasoning step. This technique provides dense feedback for policy optimization, addressing both sparse rewards and training instability. Further, some research focuses on adjusting the focus of RL algorithms to optimize exploration in targeted aspects. Specifically, CPL [639], cDPO [364], and Focused-DPO [831] enhance exploration in Long CoT by prioritizing critical or error-prone areas through preference optimization, improving accuracy in those regions. Bartoldson et al. [37] further adjusts the replay strategy of the training data, aiming to optimize reasoning performance. Li et al. [339] introduce Learning Impact Measurement (LIM), an automated method for evaluating and prioritizing training samples based on their alignment with model learning trajectories. This approach enables efficient resource use and scalable implementation. For instance, ThinkPO [746] uses short CoT reasoning outputs as rejected answers and longer ones as chosen answers for the same question, applying DPO to encourage prioritization of longer reasoning outputs [896].
@@
(2) Reward-model-based RL: Reward-model-based RL refers to approaches that use a reward model to guide learning and decision-making in the absence of explicit rewards [833]. Earlier, Proximal Policy Optimization (PPO) was first introduced by Schulman et al. [517], which alternates between interacting with the environment to collect data and optimizing a surrogate objective function via stochastic gradient ascent, surpassing DPO [239]. Subsequently, ReMax [350] eliminates the need for additional value models in PPOs. By incorporating variance reduction and REINFORCE [555] techniques, it reduces over four hyperparameters, resulting in lower GPU memory usage and faster training. Building on this, DeepSeekMath [524] proposes Group Relative Policy Optimization (GRPO), replacing traditional value models with improved sampling strategies, thus significantly accelerating learning and achieving performance on par with GPT-4 in mathematics. Hu [207] further refine GRPO with REINFORCE++, simplifying the algorithm and enhancing its training. Additionally, Vassoyan et al. [602] and [888] improve exploration efficiency in smaller models by modifying the KL penalty, thus enhancing performance under distribution shifts. Huang et al. [216] introduce Decoupled Value Policy Optimization (DVPO), a streamlined framework that replaces reward modeling with a pretrained global value model (GVM) and eliminates the interdependence between actor and critic. To address the high-quality demands of reward models, Cui et al. [115] propose PRIME (Process Reinforcement through IMplicit rEwards), which integrates the SFT model as a PRM within a unified reinforcement learning framework, enabling online updates through policy rollouts and outcome labels via implicit process rewards. Finally, Yi et al. [770] introduce SPPD, which employs Process Preference Learning with Dynamic Value Margin for self-training.
@@ 6.2.2 Reward Strategies
The studies explore advancements in training advanced RLLMs using rule-rewarded RL to enhance exploration strategies and reasoning accuracy. These efforts primarily focus on three types of rewards: (1) Correctness Rewarding: Correctness rewards are fundamental for guiding RLLMs toward accurate answers. Specifically, Singh et al. [534] introduce a binary reward system (positive or negative) to facilitate exploration, achieving simple yet effective performance improvements. Similarly, the DeepSeek-R1 [179] employs rule-extracted accuracy as an RL reward, scaling this approach to larger scenarios and training sizes, thereby enhancing both exploration and reasoning tasks [411, 136]. Furthermore, O1-CoderZhang et al. [854], StepCoder [128], and SWE-RL [670] address challenges in code generation by developing a test case generator, which standardizes code testing, ensuring accurate generation [708, 788]. (2) Format Rewarding: Further, format rewards are used to encourage better reasoning paradigms. Guo et al. [179] introduce this concept to effectively guide reasoning and exploration [701]. Xie et al. [701] expanded on this with a three-stage, rule-based RL approach, enabling the Qwen-7B model to learn complex multi-path exploration, which significantly improved both output format and corresponding length consistency. (3) Scaling rewarding: Moreover, scaling rewards are applied to promote longer reasoning chains and broader exploration. Recent studies [78, 462, 277] highlight the need for progressively scaled reasoning lengths to overcome the limitations of current reasoning approaches. As a result, research has focused on scaling exploration [701, 762]. However, excessive scaling can lead to inefficiency and overcomplicated reasoning [114]. Kimi-K1.5 [572], Yang et al. [747] and Arora and Zanette [19] proposed Long2Short techniques, favoring shorter, more accurate reasoning may also significantly improve efficiency and performance.
@@
It refers to a class of techniques in which RL algorithms are enhanced by leveraging additional reward models, to guide exploration and improve decision-making processes [547]. Earlier in 2021, OpenAI [113] propose a “Gen-Verifier” paradigm to train a correctness-oriented ORM and used ORM-rewarded RL to surpass SFT performance. Recently, with rapid advancements in PRM, several studies [605, 822, 408] have scaled reinforcement learning by enhancing exploration through step-level correctness rewarding [525]. Building on this, Hou et al. [206] introduce entropy rewards and dynamic regularization to further optimize the reasoning process. STeCa [617] identifies suboptimal actions during exploration by comparing step-level rewards and adjusting trajectories to improve deep reasoning. Additionally, the Kimi-K1.5 model [572] extends PRM paradigms into multimodal scenarios, achieving state-of-the-art performance in multi-modal reasoning tasks through a streamlined reinforcement learning framework.
@@ 6.3.1 Human-driven Exploration
Human-driven exploration refers to human-designed constant pipeline exploration for long-term exploration. Several studies highlight the effectiveness of prompt-based [268, 588, 167], tree-structured [884, 755, 81, 496, 438, 43] and even graph-structured [42, 584, 485, 56, 847] search frameworks, demonstrating superior performance and scalability over traditional methods across various datasets. Building on this, CodeTree [323] and Tree-of-Code [446] integrate a tree-based structure with execution and LLM feedback, utilizing multi-agents to optimize multi-stage decisions, thereby improving both strategy planning and solution refinement. Cheng et al. [99] generalize this approach with the Self-Play with Tree-Search Refinement (SPAR) strategy, which generates valid, comparable preference pairs to enhance instruction-following capabilities. Bi et al. [47] and Light et al. [357] extend tree search to a multi-tree paradigm, introducing the Forest-of-Thought framework, which incorporates multiple reasoning trees to improve exploration capabilities to solve complex tasks with greater accuracy. Furthermore, Li et al. [312] explores the integration of Python tools into Long CoT frameworks by both prompting and training, performing test-time scaling more effectively.
@@ 6.3.2 Model-driven Exploration
Recent efforts have focused on improving exploration structures during iterations for better logical quality. (1) Beam Search: Earlier, Xie et al. [703] introduced a decoding algorithm that integrates self-evaluation guidance via stochastic beam search, using it as a more reliable automatic criterion to streamline the search in the reasoning space, thereby enhancing prediction quality. Similarly, Zhu et al. [902] propose Deductive Beam Search (DBS), which combines CoT and deductive reasoning with stepwise beam search for RLLMs. (2) A* Search: On another front, Lehnert et al. [304] present Searchformer, which predicts A* algorithm dynamics to improve task performance and reduce search steps [85]. Later, Kang et al. [267] introduce the MindStar (M*) framework, which optimizes reasoning paths through beam search and Levin tree search methods, further enhancing reasoning performance. (3) MCTS Search: Building on the advantages of MCTS, a series of studies, such as Macro-o1 [867], STILL-1 [255], SRA-MCTS [710], and RFTT [833], utilize MCTS to guide more effective exploration [828, 333, 264, 253, 877, 488, 465]. Xu [714] utilizes energy function for better exploration during Long CoT. Yao et al. [753] further advance this by introducing Collective MCTS (CoMCTS), which leverages collective learning across multiple LLMs to enhance reasoning. Further, MC-NEST [499] integrates Nash Equilibrium strategies to balance exploration and exploitation, improving LLM decision-making in multi-step mathematical tasks. Additionally, CoAT [455] expands the MCTS algorithm with a dynamic correlation memory mechanism, enabling the system to dynamically store new information during inference. Despite MCTS’s benefits, it is often hindered by a large action space and inefficient search strategies, which complicate the generation of Long CoTs. To address this, Lin et al. [361] propose constraining the action space and refining the search strategy to facilitate the emergence of long CoTs. Finally, these methods have been extended to interactive environments, significantly improving success rates in automated exploration tasks [613, 283, 356, 707, 815, 463].
@@
MedThoughts-8K Medicine DeepSeek R1 Lang 8K
SYNTHETIC-1 [429] Mathematics, Code, Science DeepSeek R1 Lang 894K
@@ 7.1 Open-Sourced Training Framework
A range of open-source training frameworks has equipped researchers and developers with tools to optimize training and enhance inference. Each framework is built on distinct design principles and features. Early frameworks like SimpleRL [809] and DeepScaler [408] quickly replicated R1’s technology stack. Others, such as X-R1 [583] and TinyZero [456], emphasize delivering an intuitive “Aha moment” experience for under \$50. Open-Reasoner-Zero [209] replicated the DeepSeek-R1-zero training scheme with a 32B model and achieved a similar performance. Additionally, LLM Reasoner [186] provides tools to help researchers adapt strategies for External Exploration. Frameworks such as OpenR [624], OpenRLHF [208], OpenR1 [571], and Logic-RL [701] have enhanced the replication of Long CoT in deep reinforcement learning for text modalities. Further, DAPO [781] and VAPO [801] DAPO [781] and VAPO [801] enhance the efficiency of Long CoT RL training by incorporating more detailed and fine-grained training strategies. R1-V [74], R1-Multimodal-Journey [522], VL-Thinking [67], VLM-R1 [526], Open-R1-Multimodal [289], and Video-R1 [582] have extended the R1 framework to multimodal settings, enabling cross-modal R1-like reinforcement learning-based training. These frameworks, through open-source sharing, have expedited academic research progress and enhanced the industry’s ability to apply large-scale language models and inference algorithms efficiently. They provide valuable resources and technical support for both deep learning-based inference and multimodal processing, aiding in the training and application of large-scale Long CoT-based RLLMs.
@@ 8.1 Multimodal Long CoT
Recent discussions have focused on extending reasoning chains to multimodal contexts in the areas of Long CoT and multimodal reasoning [491, 423, 705, 689, 818, 802, 395, 194, 717, 419]. Zhang et al. [858] introduce multimodal chain-of-thought (MMCoT), while M3CoT [79] extends this with complex MMCoT, similar to Long CoT, and provides an evaluation benchmark. This work suggests that mimicking human Long CoT offers an effective solution [221, 188]. Multimodal Long CoT can be categorized into three main approaches: (1) Multimodal Long CoT Prompting: Eariler, Chen et al. [79] demonstrate that the basic description-then-reasoning prompt fails in Long CoT scenarios. To fill this gap, Li et al. [346] improve Vision RLLMs by enabling detailed, context-aware descriptions through an iterative self-refinement loop, allowing interactive reasoning for more accurate predictions without additional training. Dong et al. [126] incorporate multi-agent interaction during prompting, further scaling the reasoning length and achieving better accuracy. Furthermore, FaST [548] uses a switch adapter to select between Long CoT and direct answer modes, resulting in enhanced performance. (2) Multimodal Long CoT Imitation: Recent models such as LLaVA-CoT [713] and Virgo [133] employ data distillation to enable the imitation of Long CoT processes, addressing more complex problem-solving tasks [585]. Additionally, AtomThink [696] offers a Long CoT annotation engine that generates high-quality CoT annotations, mitigating the issue of insufficient visual mathematical data. Wei et al. [667] further extend Long CoT paradigms by incorporating more tokens during perception, improving geometric reasoning. (3) Reward Model-Based Multimodal Long CoT Exploration: Recent research employs reward or value models to enhance inference test-time scaling in both exploration and training phases. This includes model decoding [387, 52, 709] and RL training [696, 643, 815, 611, 228, 474, 558, 391], as well as the diffusion process [414, 774], all contributing to improved visual reasoning and comprehension.
The primary challenges in multimodal Long CoT are: (1) Incorporating Multimodal Reasonings: Enabling RLLMs to assist reasoning by generating [102, 182, 314] or grounding [682] visual content holds promise for improving complex spatial reasoning tasks, particularly when logic cannot be easily conveyed through text alone [314]. (2) Extending Longer Reasoning Processes: While current models focus on imitating Long CoT, there remains a lack of exploration into how multimodal inference test-time scaling can be achieved through methods like RL or MCTS [680, 241], presenting an interesting avenue for future research.

### v4

@@ 6.2 Internal Exploration
As noted in Chu et al. [131], Shen et al. [634], and Yang et al. [903], SFT serves as a memory process, while RL enhances generalization [340]. Specifically, SFT stabilizes the model’s output format, whereas RL improves its generalization capacity, which can increase learning efficiency by up to eight times in tasks such as mathematical reasoning [623]. Consequently, as shown in Figure 10, leading research emphasizes the role of RL and reward strategies in enhancing the exploration capabilities of LLMs without external assistance. The performance comparison is presented in Table 5.
@@ 6.2.1 RL Strategies
(1) Reward-free RL: The first series of work focuses on RL optimization algorithms. Additionally, OREO [745] propose an offline RL method that optimizes the soft Bellman equation, improving credit assignment for multi-step reasoning tasks and outperforming existing approaches in fields like mathematics and agent control. Liu et al. [453] propose Direct Advantage Policy Optimization, a novel offline RL method that leverages a separately trained critic to evaluate the accuracy of each reasoning step. This technique provides dense feedback for policy optimization, addressing both sparse rewards and training instability. Further, some research focuses on adjusting the focus of RL algorithms to optimize exploration in targeted aspects. Specifically, CPL [771], cDPO [434], and Focused-DPO [1004] enhance exploration in Long CoT by prioritizing critical or error-prone areas through preference optimization, improving accuracy in those regions. Bartoldson et al. [39] further adjusts the replay strategy of the training data, aiming to optimize reasoning performance. Li et al. [399] introduce Learning Impact Measurement (LIM), an automated method for evaluating and prioritizing training samples based on their alignment with model learning trajectories. This approach enables efficient resource use and scalable implementation. For instance, ThinkPO [907] uses short CoT reasoning outputs as rejected answers and longer ones as chosen answers for the same question, applying DPO to encourage prioritization of longer reasoning outputs [1089].
@@
(2) Reward-based RL: Reward-model-based RL refers to approaches that use a reward model or a verifier to guide learning and decision-making in the absence of explicit rewards [1007, 164, 622, 262, 795, 815]. Earlier, Proximal Policy Optimization (PPO) was first introduced by Schulman et al. [621], which alternates between interacting with the environment to collect data and optimizing a surrogate objective function via stochastic gradient ascent, surpassing DPO [288]. Subsequently, ReMax [415] eliminates the need for additional value models in PPOs. By incorporating variance reduction and REINFORCE [676] techniques, it reduces over four hyperparameters, resulting in lower GPU memory usage and faster training. Building on this, DeepSeekMath [631] proposes Group Relative Policy Optimization (GRPO), replacing traditional value models with improved sampling strategies, thus significantly accelerating learning and achieving performance on par with GPT-4 in mathematics. Hu [248] and Liu et al. [476] further refine GRPO with REINFORCE++ and Dr. GRPO, respectively, simplifying the algorithm and enhancing its training. Additionally, Vassoyan et al. [725] and [1079] improve exploration efficiency in smaller models by modifying the KL penalty, thus enhancing performance under distribution shifts. Huang et al. [260] introduce Decoupled Value Policy Optimization (DVPO), a streamlined framework that replaces reward modeling with a pretrained global value model (GVM) and eliminates the interdependence between actor and critic. To address the high-quality demands of reward models, Cui et al. [137] propose PRIME (Process Reinforcement through IMplicit rEwards), which integrates the SFT model as a PRM within a unified reinforcement learning framework, enabling online updates through policy rollouts and outcome labels via implicit process rewards. Finally, Yi et al. [934] introduce SPPD, which employs Process Preference Learning with Dynamic Value Margin for self-training. Liang et al. [416] introduce Self-aware Weakness-driven Problem Synthesis, a reinforcement-learning method that generates challenges tailored to an RLLM’s specific weaknesses [830, 171]. By concentrating training on its most difficult aspects, the model achieves more focused and effective reasoning improvements [653]. Wang et al. [775] introduce ROLL, a method designed to support R1-level large-scale training of RLLMs, enabling the efficient exploration and optimization of reasoning paths within the Mixture-of-Experts (MOE) structure [759]. Fu et al. [176] introduce AReaL, a large-scale asynchronous reinforcement learning system for language reasoning, which enhances the efficiency and effectiveness of training RLLMs. Ma et al. [503] introduce a novel interleaved SFT and RL method that focuses on solving the hardest questions when RL can not sample any correct answer, enabling RLLMs to learn from their mistakes and improve their reasoning capabilities.
@@ 6.2.2 Reward Strategies
The studies explore advancements in training advanced RLLMs using rule-rewarded RL to enhance exploration strategies and reasoning accuracy [279]. These efforts primarily focus on three types of rewards: (1) Correctness Rewarding: Correctness rewards are fundamental for guiding RLLMs toward accurate answers. Specifically, Singh et al. [647] introduce a binary reward system (positive or negative) to facilitate exploration, achieving simple yet effective performance improvements. Similarly, the DeepSeek-R1 [213] employs rule-extracted accuracy as an RL reward, scaling this approach to larger scenarios and training sizes, thereby enhancing both exploration and reasoning tasks [499, 160]. Furthermore, O1-CoderZhang et al. [1035], StepCoder [151], and SWE-RL [809] address challenges in code generation by developing a test case generator, which standardizes code testing, ensuring accurate generation [859, 956]. (2) Format Rewarding: Further, format rewards are used to encourage better reasoning paradigms. Guo et al. [213] introduce this concept to effectively guide reasoning and exploration [852]. Xie et al. [852] expanded on this with a three-stage, rule-based RL approach, enabling the Qwen-7B model to learn complex multi-path exploration, which significantly improved both output format and corresponding length consistency. (3) Scaling rewarding: Moreover, scaling rewards are applied to promote longer reasoning chains and broader exploration. Recent studies [85, 557, 330] highlight the need for progressively scaled reasoning lengths to overcome the limitations of current reasoning approaches. As a result, research has focused on scaling exploration [852, 926]. However, excessive scaling can lead to inefficiency and overcomplicated reasoning [136]. Kimi-K1.5 [694], Yang et al. [908] and Arora and Zanette [21] proposed Long2Short techniques, favoring shorter, more accurate reasoning may also significantly improve efficiency and performance.
@@
It refers to a class of techniques in which RL algorithms are enhanced by leveraging additional reward models, to guide exploration and improve decision-making processes [665]. Earlier in 2021, OpenAI [135] propose a “Gen-Verifier” paradigm to train a correctness-oriented ORM and used ORM-rewarded RL to surpass SFT performance. Recently, with rapid advancements in PRM, several studies [728, 994, 495] have scaled reinforcement learning by enhancing exploration through step-level correctness rewarding [632]. Building on this, Hou et al. [247] introduce entropy rewards and dynamic regularization to further optimize the reasoning process [111]. STeCa [741] identifies suboptimal actions during exploration by comparing step-level rewards and adjusting trajectories to improve deep reasoning. Additionally, the Kimi-K1.5 model [694] extends PRM paradigms into multimodal scenarios, achieving state-of-the-art performance in multi-modal reasoning tasks through a streamlined reinforcement learning framework.
@@ 6.3.1 Human-driven Exploration
Human-driven exploration refers to human-designed constant pipeline exploration for long-term exploration [456, 401]. Several studies highlight the effectiveness of prompt-based [320, 710, 199, 217, 833, 595, 530, 1026, 639], tree-structured [1075, 919, 90, 599, 531, 45, 230] and even graph-structured [44, 706, 584, 60, 1027, 1041] search frameworks, demonstrating superior performance and scalability over traditional methods across various datasets. Building on this, CodeTree [380] and Tree-of-Code [539] integrate a tree-based structure with execution and LLM feedback, utilizing multi-agents to optimize multi-stage decisions, thereby improving both strategy planning and solution refinement [684]. Cheng et al. [113] generalize this approach with the Self-Play with Tree-Search Refinement (SPAR) strategy, which generates valid, comparable preference pairs to enhance instruction-following capabilities. Bi et al. [50] and Light et al. [425] extend tree search to a multi-tree paradigm, introducing the Forest-of-Thought framework, which incorporates multiple reasoning trees to improve exploration capabilities to solve complex tasks with greater accuracy. Furthermore, Li et al. [368] explores the integration of Python tools into Long CoT frameworks by both prompting and training, performing test-time scaling more effectively.
@@ 6.3.2 Model-driven Exploration
Recent efforts have focused on improving exploration structures during iterations for better logical quality. (1) Beam Search: Earlier, Xie et al. [854] introduced a decoding algorithm that integrates self-evaluation guidance via stochastic beam search, using it as a more reliable automatic criterion to streamline the search in the reasoning space, thereby enhancing prediction quality [530]. Similarly, Zhu et al. [1099] propose Deductive Beam Search (DBS), which combines CoT and deductive reasoning with stepwise beam search for RLLMs. (2) A* Search: On another front, Lehnert et al. [359] present Searchformer, which predicts A* algorithm dynamics to improve task performance and reduce search steps [96]. Later, Kang et al. [319] introduce the MindStar (M*) framework, which optimizes reasoning paths through beam search and Levin tree search methods, further enhancing reasoning performance. (3) MCTS Search: Building on the advantages of MCTS, a series of studies, such as Macro-o1 [1053], STILL-1 [304], SRA-MCTS [862], and RFTT [1007], utilize MCTS to guide more effective exploration [1001, 391, 316, 302, 1068, 587, 560, 429]. Xu [867] utilizes energy function for better exploration during Long CoT. Yao et al. [917] further advance this by introducing Collective MCTS (CoMCTS), which leverages collective learning across multiple LLMs to enhance reasoning. Further, MC-NEST [603] integrates Nash Equilibrium strategies to balance exploration and exploitation, improving LLM decision-making in multi-step mathematical tasks [905, 1047]. Additionally, CoAT [549] expands the MCTS algorithm with a dynamic correlation memory mechanism, enabling the system to dynamically store new information during inference. Despite MCTS’s benefits, it is often hindered by a large action space and inefficient search strategies, which complicate the generation of Long CoTs. To address this, Lin et al. [430] propose constraining the action space and refining the search strategy to facilitate the emergence of Long CoTs. Finally, these methods have been extended to interactive environments, significantly improving success rates in automated exploration tasks [737, 336, 424, 858, 985, 558, 764, 442].
@@
LiveCodeBench (History) [291] Code Human Lang 0.9K
LeetCode [844] Code Human Lang 2K
ARC-AGI [126] Logic Puzzle Human Synthesis Lang 0.4K
ARC-AGI-2 [127] Logic Puzzle Human Synthesis Lang 1K
BARC [393] Logic Puzzle Human Synthesis Lang 3.4K
Code I/O (PyEdu) [381] Code Execution Simulation Human Synthesis Lang 227K
HiTab [118] Tabular Human Lang 7.5K
MultiHierTT [381] Code Execution Simulation Human Synthesis Lang 7.8K
@@
MedThoughts-8K [804] Medicine DeepSeek R1 Lang 8K
PrimeIntellect [520] Code DeepSeek R1 Lang 16.3K
@@
OR1 [228] Mathematics, Code, General Human Question + Rule Validation Lang 105K
DeepScaler [495] Mathematics Human Question + Rule Validation Lang 40.3
DAPO [948] Mathematics Human Question + Rule Validation Lang 17K
TACO-Verified [382] Code Human + Rule Validation Lang 0.9K
WebInstruct-Verifed [508] Science, General Web Crawling + Rule & LLM Validation Lang 232K
Guru92K [119] Mathematics, Code, Puzzle, General Unified + Rule Validation Lang 92K
@@ 7.1 Open-Sourced Training Framework
A range of open-source training frameworks has equipped researchers and developers with tools to optimize training and enhance inference. Each framework is built on distinct design principles and features. Early frameworks like SimpleRL [979] and DeepScaler [495] quickly replicated R1’s technology stack. Others, such as X-R1 [705] and TinyZero [550], emphasize delivering an intuitive “Aha moment” experience for under \$50. Open-Reasoner-Zero [250] replicated the DeepSeek-R1-zero training scheme with a 32B model and achieved a similar performance. Additionally, LLM Reasoner [221] provides tools to help researchers adapt strategies for External Exploration. Frameworks such as OpenR [749], OpenRLHF [249], OpenR1 [693], and Logic-RL [852] have enhanced the replication of Long CoT in deep reinforcement learning for text modalities. Further, DAPO [948] and VAPO [972] enhance the efficiency of Long CoT RL training by incorporating more detailed and fine-grained training strategies. R1-V [81], R1-Multimodal-Journey [629], VL-Thinking [74], VLM-R1 [633], Open-R1-Multimodal [342], and Video-R1 [704] have extended the R1 framework to multimodal settings, enabling cross-modal R1-like reinforcement learning-based training. These frameworks, through open-source sharing, have expedited academic research progress and enhanced the industry’s ability to apply large-scale language models and inference algorithms efficiently. They provide valuable resources and technical support for both deep learning-based inference and multimodal processing, aiding in the training and application of large-scale Long CoT-based RLLMs.
@@ 8.1 Multimodal Long CoT
Recent discussions have focused on extending reasoning chains to multimodal contexts in the areas of Long CoT and multimodal reasoning [592, 514, 856, 836, 988, 973, 478, 232, 870, 510, 407, 812, 1055]. Zhang et al. [1040] introduce multimodal chain-of-thought (MMCoT), while M3CoT [86] extends this with complex MMCoT, similar to Long CoT, and provides an evaluation benchmark. This work suggests that mimicking human Long CoT offers an effective solution [267, 223, 992]. Multimodal Long CoT can be categorized into three main approaches: (1) Multimodal Long CoT Prompting: Eariler, Chen et al. [86] demonstrate that the basic description-then-reasoning prompt fails in Long CoT scenarios. To fill this gap, a series of work focus on optimizing the multimodal Long CoT capabilities [529, 1062]. For example, Li et al. [410] improve Vision RLLMs by enabling detailed, context-aware descriptions through an iterative self-refinement loop, allowing interactive reasoning for more accurate predictions without additional training. Dong et al. [149] incorporate multi-agent interaction during prompting, further scaling the reasoning length and achieving better accuracy. Furthermore, FaST [667] uses a switch adapter to select between Long CoT and direct answer modes, resulting in enhanced performance. (2) Multimodal Long CoT Imitation: Recent models such as LLaVA-CoT [866] and Virgo [156] employ data distillation to enable the imitation of Long CoT processes, addressing more complex problem-solving tasks [707, 92, 637]. Additionally, AtomThink [845] offers a Long CoT annotation engine that generates high-quality CoT annotations, mitigating the issue of insufficient visual mathematical data. Wei et al. [805] further extend Long CoT paradigms by incorporating more tokens during perception, improving geometric reasoning. (3) Reward Model-Based Multimodal Long CoT Exploration: Recent research employs reward or value models to enhance inference test-time scaling in both exploration and training phases. This includes model decoding [466, 56, 860, 885] and RL training [845, 776, 985, 734, 276, 571, 679, 474, 414], as well as the diffusion process [504, 939, 850], all contributing to improved visual reasoning and comprehension.
The primary challenges in multimodal Long CoT are: (1) Incorporating Multimodal Reasonings: Enabling RLLMs to assist reasoning by generating [120, 216, 370, 122] or grounding [824] visual content holds promise for improving complex spatial reasoning tasks, particularly when logic cannot be easily conveyed through text alone [121, 666, 91, 878]. (2) Extending Longer Reasoning Processes: While current models focus on imitating Long CoT, there remains a lack of exploration into how multimodal inference-time scaling can be achieved through methods like RL or MCTS [822, 290], presenting an interesting avenue for future research [468].